OpenStudy (anonymous):

The volume of the cage must be 64m^3. the width to be 4 meters and the height of the cage to be 6 meters less than the length if the other side of the cage. L= and H=

OpenStudy (anonymous):

L*W*H=64 we are told that W=4, so L*4*H=64 (we will come back to this to find H) the height is 6 less than the Length translates to H=L-6, so L*4*(L-6)=64 which simplifies to 4L^2-24L=64 subtract 64 from both sides 4L^2-24L-64=0 dividing by 4 4L^2-24L-64=0 L^2-6L-16=0 factoring (L-8)(L+2)=0 setting each factor to 0 and solving L=8 or L=-2 so, L=8 To get H, substitute this result into the equation above L*4*H=64: 8*4*H=64 32H=64 H=2 Answer: L=8; H=2 Check: L*W*H=64 8*4*2=32*2=64

OpenStudy (anonymous):

Thank you so much I've been stuck on this for awhile. It turns out the way I was going about it was wrong.

OpenStudy (anonymous):

You are welcome!